# CHG Shapley: Efficient Data Valuation and Selection towards Trustworthy Machine Learning

### v2

@@ Abstract
Understanding the decision-making process of machine learning models is crucial for ensuring trustworthy machine learning. Data Shapley, a landmark study on data valuation, advances this understanding by assessing the contribution of each datum to model accuracy. However, the resource-intensive and time-consuming nature of multiple model retraining poses challenges for applying Data Shapley to large datasets. To address this, we propose the CHG (Conduct of Hardness and Gradient) score, which approximates the utility of each data subset on model accuracy during a single model training. By deriving the closed-form expression of the Shapley value for each data point under the CHG score utility function, we reduce the computational complexity to the equivalent of a single model retraining, an exponential improvement over existing methods. Additionally, we employ CHG Shapley for real-time data selection, demonstrating its effectiveness in identifying high-value and noisy data. CHG Shapley facilitates trustworthy model training through efficient data valuation, introducing a novel data-centric perspective on trustworthy machine learning.
@@ 1 Introduction
The central problem of trustworthy machine learning is explaining the decision-making process of models to enhance the transparency of data-driven algorithms. However, the high complexity of machine learning model training and inference processes obscures an intuitive understanding of their internal mechanisms. Approaching trustworthy machine learning from a data-centric perspective (Liu et al., 2023) offers a new perspective for research. For trustworthy model inference, a representative algorithm is the SHAP (Lundberg and Lee, 2017), which quantitatively attributes model outputs to input features, clarifying which features influence specific results the most. SHAP and its variants(Kwon and Zou, 2022b) are widely applied in data analysis and healthcare. For trustworthy model training, the Data Shapley algorithm (Ghorbani and Zou, 2019) stands out. It quantitatively attributes a model’s performance to each training data point, identifying valuable data that improves performance and noisy data that degrades it. Data valuation not only informs us about the extent to which each training data impacts model performance but also underpins subsequent tasks such as data selection, data acquisition, data cleaning, etc., and supports the development of data markets (Mazumder et al., 2023).
@@
The impressive effectiveness of both SHAP and Data Shapley algorithms is rooted in the Shapley value (Shapley, 1953). The unique feature of the Shapley value lies in its ability to accurately and fairly allocate contributions to each factor in decision-making processes where multiple factors interact with each other. However, the exact computation of the Shapley value is $\mathcal{O}(2^{n})$, where $n$ is the number of participants. Consequently, even when employing estimation methods akin to linear least squares regression (Lundberg and Lee, 2017) or Monte Carlo (Ghorbani and Zou, 2019), the efficiency of SHAP and Data Shapley algorithms sometimes fails to meet requirements. For example, the most efficient known data valuation method is KNNShapley (Jia et al., 2019) designed for the k-nearest neighbors algorithm, requiring $\mathcal{O}(\sqrt{n}\log(n)^{2})$ times of model retraining, where $n$ is the number of training data.
@@
In this paper, we focus on the efficiency problem of data valuation on large-scale datasets. Instead of investigating more efficient and robust algorithms to approximate the Shapley value like in (Lundberg and Lee, 2017; Wang and Jia, 2023; Wang et al., 2024b), we concentrate on estimating the utility function. The underlying idea is that if the accuracy estimation function of a model trained on a subset can be analytically expressed through some features of the subset, then it becomes possible to directly derive an analytic expression for the Shapley value of each data point.
Our contributions are summarized as follows:
@@
Efficient and Large-Scale Data Valuation Method: We introduce the CHG (Conduct of Hardness and Gradient) score, which assesses the influence of a data subset on model accuracy. By deriving the analytical expression for the Shapley value of each data point under this utility function, we significantly improve upon the $\mathcal{O}(n^{2}\log n)$ time complexity of Data Shapley. Using our proposed CHG Shapley method, valuing the full training dataset takes at most twice the time of conventional training, equivalent to one additional model retraining.
@@
Real-Time Selection of Training Data in Large-Scale Datasets Based on Data Valuation: We employ CHG Shapley for real-time data selection. Experiments involving data-efficient training and model training under noise labels demonstrate the effectiveness of CHG Shapley in identifying high-value and noisy data.
@@
A New Data-Centric Perspective on Trustworthy Machine Learning: CHG Shapley is a parameter-free method, deriving its benefits solely from a deeper understanding of the data. We believe that approaches rooted in data valuation have the potential to enhance understanding of model mechanisms, thereby advancing trustworthy machine learning.
@@ 2 Background and Related Works
Shapley value. Supposing some players cooperate to finish a task, we denote the player set as $N$ and the utility function defined on every subset of $N$ is denoted as $U$, then the Shapley value (Shapley, 1953) is the unique solution concept that satisfies the following four axioms:
@@
A milestone in data valuation research is the Data Shapley algorithm developed by Ghorbani and Zou (2019). Data Shapley builds upon the Shapley value formula by treating all data in the training dataset as players in a cooperative game of "training the model together". $U(S)$ represents the accuracy achieved by training the model on dataset subset $S$. Consequently, the Shapley value computed for each data point serves as an indication of its contribution to model performance. Data points with lower values are considered more likely to be harmful, while those with higher values are deemed to be lacking in the model’s representation. Removing low-value data points or acquiring more high-value data points can lead to improved model accuracy, a hypothesis supported by related experiments. Following Data Shapley, several subsequent have been proposed.
KNN Shapley(Jia et al., 2019) focus on the Shapley value of the k-nearest neighbors algorithm, proposing the closed-form expression of the Shapley value of the KNN scenario, notably improving computational efficiency to $\mathcal{O}(\sqrt{n}\log(n)^{2})$.
Beta Shapley (Kwon and Zou, 2022a) was developed to address an issue identified in Data Shapley regarding the uniform weight allocation for sets with different cardinalities. By relaxing the efficiency axiom in the Shapley value and utilizing a beta distribution to characterize the weights of sets of different cardinalities, Beta Shapley performs better than Data Shapley in tasks such as detecting mislabeled training data.
Data Banzhaf (Wang and Jia, 2023) aims to mitigate the noise introduced by stochastic gradient algorithms on model accuracy (i.e., utility function) during the model training process. Banzhaf value has also been demonstrated to be the most robust solution concept among all semi-values (i.e., Shapley values without Efficiency restrictions) in the presence of noise in the utility function. Therefore, Data Banzhaf is well-suited for evaluating training data in deep learning scenarios.
Additionally, there has been a surge in work on machine learning systems that utilize Shapley values to compute data values (Karlaš et al., 2024). All these studies utilize the Shapley value or its variants as their theoretical foundation because it is widely acknowledged in cooperative game theory as the fairest method for allocating contributions (Algaba et al., 2019).
@@
To provide a more detailed assessment of various data valuation algorithms, Jiang et al. (2023) introduced the OpenDataVal benchmark library. This benchmark library abstracts away details such as model training, allowing researchers to focus solely on algorithm design for data valuation. We also use this benchmark to evaluate our proposed data valuation algorithm.
@@ 3 Efficient Data valuation
"What’s the value of your data?" Instead of directly assessing the contribution of each data point to model accuracy, we initially explore another pertinent question: "Which data holds greater value in the training process?" In the realm of data selection, numerous significant studies offer varying perspectives. Some focus on the utility of individual data, employing methods such as influence functions (Koh and Liang, 2017; Yang et al., 2023; Chhabra et al., 2024), EL2N score (Paul et al., 2021), InfoBatch (Qin et al., 2024), and worst-case training methods (Huang et al., 2022). These approaches utilize data information and model parameters to calculate a score (e.g., loss) for data selection. Conversely, another approach emphasizes the utility of data subsets (Killamsetty et al., 2021a, b; Zhao et al., 2021), aiming to identify representative subsets of the original dataset that can achieve comparable model accuracy, often utilizing corresponding gradient subsets. We reconcile these two perspectives and introduce the CHG score as a measure of data subset accuracy.
@@ 3.1 Preliminaries
Let $N=\{1,\dots,n\}$ denotes a labeled training set. The objective is to assign a scalar to each training data point based on its contribution to the model’s performance. A utility function $U:2^{N}\rightarrow R$, mapping from any subset $S$ of the training set (including the empty set) to the performance of the model trained on the subset (denoted as $U(S)$), is usually used for data valuation. In Data Shaply(Ghorbani and Zou, 2019), the accuracy of the model on a hold-out test set is used as $U(S)$.
@@ 3.2 CHG Score as a Utility Function for Subset Model Performance
The opacity of $U(S)$, or the fact that $U(S)$ can only be obtained after training a model on $S$, renders marginal contribution-based data valuation algorithms such as Data Shapley (Ghorbani and Zou, 2019), Beta Shapley (Kwon and Zou, 2022a), Data Banzhaf (Wang and Jia, 2023), AME (Lin et al., 2022), and LOO (leave-one-out error) time-consuming and computationally intensive.
On the contrary, if we can determine the model’s accuracy on $S$ without training on it, or even derive the expression for $U(S)$ in terms of $S$, then we can significantly reduce the model training time and expedite the calculation of each data’s value. To achieve this, we first investigate the impact of training the model using a subset of data $S$.
Here we focus on the model training process. Assuming that in the current iteration, the model parameters are denoted by $\theta$. We use $f(\theta):=\frac{1}{N}\sum_{i\in N}f(i;\theta)$ to represent the loss function of the neural network across the entire training dataset. A smaller loss indicates higher model accuracy on the training dataset.
Assuming at this point, no overfitting has occurred, i.e., the model’s accuracy on the training and test datasets is comparable, we can reasonably utilize the training dataset’s loss to estimate the model’s accuracy on the test dataset. If we update the parameters by training on the subset $S$ using gradient descent, then the next update direction of the parameters is $x:=\frac{1}{|S|}\sum_{i\in S}x_{i}$, where $x_{i}$ represents the gradient of the $i$-th data point. After one step of gradient descent, $\theta$ becomes $\theta-\eta x$. Subsequently, based on the following lemma, $\|\nabla f(\theta)\|_{2}^{2}-\|\nabla f(\theta)-x\|_{2}^{2}$ can be considered as a metric for evaluating the model’s performance.
@@
Let $f$ be any differentiable function with $L$-Lipschitz continuous gradient, $\theta_{x}=\theta-\eta x$, and $\eta=1/L$ [^1] This is a commonly used assignment method, for example, in proving the convergence of gradient descent for non-convex functions to local minima (Nesterov, 2018), $\eta$ was set as $1/L$, then we have:
$f(\theta_{x})\leq f(\theta)+\langle\nabla f(\theta),\theta_{x}-\theta\rangle+\frac{L}{2}\|\theta_{x}-\theta\|_{2}^{2}=f(\theta)-\frac{\eta}{2}(\|\nabla f(\theta)||_{2}^{2}-\|\nabla f(\theta)-x||_{2}^{2})$ (2)
Proof: The first inequality is due to the definition of $L$-Lipschitz continuity.
When using traditional empirical risk minimization, measuring the model’s performance using the gradients of the training set on the training process is acceptable. However, inspired by the principles of worst-case training method (Huang et al., 2022) and InfoBatch (Qin et al., 2024), we introduce a method to incorporate the difficulty of data points into the model optimization process.
Specifically, our optimization objective is to minimize $\frac{1}{N}\sum_{i\in N}h_{i}f(i;\theta)$ rather than $\frac{1}{N}\sum_{i\in N}f(i;\theta)$ (i.e., empirical risk minimization), where $h_{i}$ is equal to $f(i;\theta)$ but does not participate in model backpropagation. $h_{i}$ can be viewed as the hardness of learning this data point. Generally, when there is no label noise, focusing on hard-to-learn data points can help improve the model’s generalization ability (Huang et al., 2022) and may reduce training time (Qin et al., 2024). Then the gradient of $i$-th data changes from $x_{i}$ to $h_{i}x_{i}$, and then we define the Conduct of Hardness and Gradient score (CHG score) of a data subset $S$ as $\|\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)\|_{2}^{2}-\|\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)-\frac{1}{S}\sum_{i\in S}l_{i}\nabla f(i;\theta)\|_{2}^{2}$. We use this score to measure the influence of data subsets on model accuracy. We also investigate the performance of the Hardness score ( $\frac{1}{S}\sum_{i\in S}l_{i}$) and Gradient Score ( $\|\frac{1}{N}\sum_{i\in N}\nabla f(i;\theta)\|_{2}^{2}-\|\frac{1}{N}\sum_{i\in N}\nabla f(i;\theta)-\frac{1}{S}\sum_{i\in S}\nabla f(i;\theta)\|_{2}^{2}$) in Table 2, see Hardness Shapley and Gradient Shapley.
@@
Data Shapley (Ghorbani and Zou, 2019) Marginal contribution $\mathcal{O}(n^{2}\log n)$ + + ++ ++
@@
Beta Shapley (Kwon and Zou, 2022a) Marginal contribution $\mathcal{O}(n^{2}\log n)$ + + ++ ++
Data Banzhaf (Wang and Jia, 2023) Marginal contribution $\mathcal{O}(n\log n)$ + +
@@
Influence Function (Koh and Liang, 2017) Gradient NA - - + +
@@
Data-OOB (Kwon and Zou, 2023) Out-of-bag estimate NA ++ + - ++
CHG Shapley (this paper) Marginal contribution, $1$ ++ + ++ NA
@@
Table 1: A taxonomy of data valuation algorithms, mainly abstracted from (Jiang et al., 2023). The symbol “NA” means the method is not based on model training, then “No Answer”. The symbols ‘- / + / ++’ indicate that a corresponding data valuation method achieves a ‘similar or worse / better / much better’ performance than a random baseline, respectively.
3.3 CHG Shapley
After obtaining the utility function (CHG score), which approximates the influence of data subsets on model accuracy using gradient and loss information during model training, we can derive the analytical expression for the Shapley value of each data point under this utility function. This enables efficient computation of the value of each data point in large-scale datasets.

### v3

@@ Abstract
Understanding the decision-making process of machine learning models is crucial for ensuring trustworthy machine learning. Data Shapley, a landmark study on data valuation, advances this understanding by assessing the contribution of each datum to model performance. However, the resource-intensive and time-consuming nature of multiple model retraining poses challenges for applying Data Shapley to large datasets. To address this, we propose the CHG (compound of Hardness and Gradient) utility function, which approximates the utility of each data subset on model performance in every training epoch. By deriving the closed-form Shapley value for each data point using the CHG utility function, we reduce the computational complexity to that of a single model retraining, achieving a quadratic improvement over existing marginal contribution-based methods. We further leverage CHG Shapley for real-time data selection, conducting experiments across three settings: standard datasets, label noise datasets, and class imbalance datasets. These experiments demonstrate its effectiveness in identifying high-value and noisy data. By enabling efficient data valuation, CHG Shapley promotes trustworthy model training through a novel data-centric perspective. Our codes are available at https://github.com/caihuaiguang/CHG-Shapley-for-Data-Valuation and https://github.com/caihuaiguang/CHG-Shapley-for-Data-Selection.
@@ 1 Introduction
The central problem of trustworthy machine learning is explaining the decision-making process of models to enhance the transparency of data-driven algorithms. However, the high complexity of machine learning model training and inference processes obscures an intuitive understanding of their internal mechanisms. Approaching trustworthy machine learning from a data-centric perspective (Liu et al., 2023) offers a new perspective for research. For trustworthy model inference, a representative algorithm is the SHAP (Lundberg & Lee, 2017), which quantitatively attributes model outputs to input features, clarifying which features influence specific results the most. SHAP and its variants(Kwon & Zou, 2022b) are widely applied in data analysis and healthcare. For trustworthy model training, the Data Shapley algorithm (Ghorbani & Zou, 2019) stands out. It quantitatively attributes a model’s performance to each training data point, identifying valuable data that improves performance and noisy data that degrades it. Data valuation reveals how much each training sample affects model performance, serving as a foundation for tasks like data selection, acquisition, and cleaning, while also facilitating the creation of data markets (Mazumder et al., 2023).
@@
The impressive effectiveness of both SHAP and Data Shapley algorithms is rooted in the Shapley value (Shapley, 1953). The unique feature of the Shapley value lies in its ability to accurately and fairly allocate contributions to each factor in decision-making processes where multiple factors interact with each other. However, the exact computation of the Shapley value is $\mathcal{O}(2^{n})$, where $n$ is the number of factors. Even with estimation techniques such as linear least squares regression (Lundberg & Lee, 2017) or Monte Carlo methods (Ghorbani & Zou, 2019), the efficiency of SHAP and Data Shapley algorithms struggles to scale with high-dimensional inputs or large datasets.
@@
In this paper, we focus on the efficiency problem of data valuation on large-scale datasets. Instead of investigating more efficient and robust algorithms to approximate the Shapley value as in (Lundberg & Lee, 2017; Wang et al., 2024b), we concentrate on estimating the utility function. The key intuition is that if the performance of a model trained on a data subset can be expressed analytically, a closed-form solution for each datum’s Shapley value could be derived, offering a more computationally efficient solution. Our contributions are summarized as follows:
@@
Efficient and Large-Scale Data Valuation Method: We introduce the CHG (Compound of Hardness and Gradient) score, which assesses the influence of a data subset on model accuracy. By deriving the analytical expression for the Shapley value of each data point under this utility function, we significantly improve upon the $\mathcal{O}(n^{2}\log n)$ time complexity of Data Shapley to a single model training run.
@@
Real-Time Training Data Selection in Large Datasets: Calculating data value was considered computationally intensive, making real-time data selection based on data value impractical. Due to the efficient computation of CHG Shapley, we further employ it for real-time data selection. Experiments conducted across three settings—standard datasets, label noise datasets, and class imbalance datasets—demonstrate CHG Shapley’s effectiveness in identifying high-value and noisy data.
@@
A New Data-Centric Perspective on Trustworthy Machine Learning: In real-world scenarios, data may not only be subject to noise and class imbalance, but it can also involve more complex mixtures of these issues. CHG Shapley is a parameter-free method that derives its advantages solely from a deeper understanding of the data. We believe that the proposed method, grounded in data valuation, has the potential to enhance trustworthy model training in complex data distributions.
@@ 2 Background and Related Works
Shapley value. The Shapley value (Shapley, 1953) is widely regarded as the fairest method for allocating contributions (Rozemberczki et al., 2022; Algaba et al., 2019). Given a set of players $N$ and a utility function $U$ defined on subsets of $N$, the Shapley value is the only solution that satisfies the axioms of dummy player, symmetry, efficiency, and linearity:
@@
A more intuitive form for the Shapley value is:
$\phi_{i}(U)=\mathbb{E}_{\pi\sim\Pi}\left[U\left(S_{\pi}^{i}\cup\{i\}\right)-U\left(S_{\pi}^{i}\right)\right].$ (2)
Here, $\pi\sim\Pi$ refers to a uniformly random permutation of the player set $N$, and $S_{\pi}^{i}$ denotes the set of players preceding player $i$ in the permutation $\pi$. This expression highlights that the Shapley value for player $i$ captures the expected marginal contribution of that player across all possible subsets. Fig. 1 illustrates the process of computing Shapley values.
Figure 1: Shapley value’s calculation.
@@
A milestone in data valuation research is the Data Shapley algorithm developed by Ghorbani & Zou (2019). Data Shapley builds upon the Shapley value formula by treating all data in the training dataset as players in a cooperative game of “training the model together”. $U(S)$ represents the accuracy achieved by training the model on dataset subset $S$. Consequently, the Shapley value computed for each data point serves as an indication of its contribution to model performance. Data points with lower values are considered more likely to be harmful, while those with higher values are deemed to be lacking in the dataset. Removing low-value or acquiring more high-value data can improve accuracy. Following Data Shapley, several subsequent have been proposed.
KNN Shapley (Jia et al., 2019) focuses on the Shapley value for k-nearest neighbors, offering a closed-form solution that improves computational efficiency to $\mathcal{O}(\sqrt{n}\log(n)^{2})$. Beta Shapley (Kwon & Zou, 2022a) addresses the uniform weighting issue in Data Shapley by relaxing the efficiency axiom and using a beta distribution, improving performance in tasks like mislabeled data detection. Data Banzhaf (Wang & Jia, 2023) reduces noise in model training by utilizing the Banzhaf value, shown to be robust in noisy environments. LAVA (Just et al., 2023) developed a proxy for validation performance, enabling the evaluation of data in a manner that is agnostic to the learning algorithms. DVRL (Kwon & Zou, 2023) and Data-OOB (Yoon et al., 2020) utilize reinforcement learning and out-of-bag estimation, respectively, to assess data values.
Several other works use gradient information to compute the value of training data. Xu et al. (2021) propose the cosine GradE Shapley value (CGSV) to measure each agent’s contribution in a collaborative machine learning scenario. CGSV approximates an agent’s Shapley value during the aggregation process by using the cosine similarity between the agent’s gradient and the aggregated gradient of all agents. Pruthi et al. (2020) introduced TracIn, a method that measures the influence of a training example on a test example by computing the inner product between their respective gradients.
@@
Finally, Jiang et al. (2023) introduced the OpenDataVal benchmark, which focuses on data valuation algorithm design by abstracting model training. We also use this benchmark to evaluate our proposed algorithm in data valuation experiments.
Data selection. There are two main streams in the data selection field. The first focuses on determining the scores of individual data points, employing methods such as influence functions (Koh & Liang, 2017; Yang et al., 2023; Chhabra et al., 2024), the EL2N score (Paul et al., 2021), and loss (i.e., hardness) metrics (e.g., InfoBatch (Qin et al., 2024), worst-case training (Huang et al., 2022)). These approaches typically use the information from a single data point and model parameters to calculate scores for that point, subsequently selecting the data points with the highest scores. Recently, Xia et al. (2024) proposed LESS, which sets the gradient mean of data within a specific downstream task as the target gradient and employs the inner product or cosine similarity to select data with the most similar gradients from a large collection of instruction datasets for model fine-tuning. Their experiments (Xia et al., 2024) show that training on just 5% of the data selected by LESS often outperforms training on the full dataset across various downstream tasks.
The second approach, including coreset methods such as GradMatch (Killamsetty et al., 2021a) and Glister (Killamsetty et al., 2021b), as well as data distillation (Zhao et al., 2021), emphasizes the utility of data subsets. These methods aim to identify a subset that effectively represents the original dataset while achieving comparable model accuracy. They typically leverage the gradients of individual training data points to form a subset whose gradient direction aligns with that of the validation set.
Our data selection method, CHG Shapley, which is grounded in data valuation principles, represents a fusion of these two streams: it leverages the utility of different training subsets to calculate a score (i.e., Shapley value) for each individual data point. We will compare our method with these two approaches in data selection experiments, demonstrating its ability to identify high-value (i.e., representative) data and low-value (i.e., noisy) data.
@@ 3.1 Preliminaries
Let $N=\{1,\dots,n\}$ denote a labeled training set. The objective is to assign a scalar value to each training data point based on its contribution to the model’s performance. A utility function $U:2^{N}\rightarrow\mathbb{R}$ maps any subset $S$ of the training set (including the empty set) to the performance of the model trained on that subset, denoted as $U(S)$. In Data Shapley (Ghorbani & Zou, 2019), the model’s accuracy on a hold-out test set is utilized as $U(S)$. Besides, similar to Data Shapley, we do not impose any distributional assumptions on the data, and independence among the data points is not required.
@@ 3.2 CHG Score as a Utility Function for Subset Model Performance
The opacity of $U(S)$, or the fact that $U(S)$ can only be obtained after training a model on $S$, renders marginal contribution-based data valuation algorithms such as Data Shapley (Ghorbani & Zou, 2019), Beta Shapley (Kwon & Zou, 2022a), Data Banzhaf (Wang & Jia, 2023), AME (Lin et al., 2022), and LOO (leave-one-out error) time-consuming and computationally intensive. On the contrary, if we can determine the model’s performance on $S$ without training on it (Wu et al., 2022; Ki et al., 2023), or even derive the expression for $U(S)$ in terms of $S$, then we may significantly reduce the model training time and expedite the calculation of each data’s value. To achieve this, we first investigate the impact of training the model using a subset of data $S$.
The objective of model training is to minimize training loss while ensuring generalization (Zhang et al., 2017) through an appropriate dataset, sufficient data quantity, model complexity, and regularization techniques, etc. Once the dataset, model architecture, regularization techniques, and other methods that ensure generalization are established, the goal with all training data is to minimize training loss as much as possible. Therefore, in our study, the utility function is defined as the degree of decrease in training loss.
Let the loss function of the neural network across the entire training dataset be represented as $f(\theta):=\frac{1}{N}\sum_{i\in N}f(i;\theta)$. When training the model on a subset $S$ using gradient descent, the next update direction for the parameters is given by $x:=\frac{1}{|S|}\sum_{i\in S}x_{i}$, where $x_{i}$ represents the gradient of the $i$-th data point. After one step of gradient descent, the parameters are updated as $\theta\leftarrow\theta-\eta x$. Based on the following lemma, the term $\|\nabla f(\theta)\|_{2}^{2}-\|\nabla f(\theta)-x\|_{2}^{2}$ can be employed as a utility function to evaluate the impacts of the subset $S$ on the training loss.
@@
Let $f$ be any differentiable function with $L$-Lipschitz continuous gradient, $\theta_{x}=\theta-\eta x$, and $\eta=1/L$ [^1] This is a commonly used assignment method, for example, in proving the convergence of gradient descent for non-convex functions to local minima (Nesterov, 2018), $\eta$ was set as $1/L$., then we have:
$f(\theta_{x})\leq f(\theta)-\frac{1}{2L}(\|\nabla f(\theta)||_{2}^{2}-\|\nabla f(\theta)-x\|_{2}^{2})$ (3)
Proof: Due to the definition of $L$-Lipschitz continuity, We have $f(\theta_{x})\leq f(\theta)+\langle\nabla f(\theta),\theta_{x}-\theta\rangle+\frac{L}{2}\|\theta_{x}-\theta\|_{2}^{2}$ Plugging in $\theta_{x}=\theta-\eta x$ to it, we get $\langle\nabla f(\theta),\theta_{x}-\theta\rangle=\langle\nabla f(\theta),-\eta x\rangle$ and $\|\theta_{x}-\theta\|_{2}^{2}=\eta^{2}\|x\|_{2}^{2}$. And with $\eta=1/L$, we then have $f(\theta)+\langle\nabla f(\theta),\theta_{x}-\theta\rangle+\frac{L}{2}\|\theta_{x}-\theta\|_{2}^{2}=f(\theta)+\langle\nabla f(\theta),-\eta x\rangle+\frac{L}{2}\eta^{2}\|x\|_{2}^{2}=f(\theta)-\frac{1}{2L}(2\langle\nabla f(\theta),x\rangle-\|x\|_{2}^{2})=f(\theta)-\frac{1}{2L}(\|\nabla f(\theta)||_{2}^{2}-\|\nabla f(\theta)-x\|_{2}^{2})$.
Lemma 1 is particularly important to our whole method. It suggests that the Euclidean distance of two gradients is not just a similarity metric like inner product or cosine, but also gives an upper bound of the training loss, which then serves as the utility function of any data subset to calculate the Shapley value of each single data.
In general, when there is no label noise, focusing on hard-to-learn data points can enhance the model’s generalization ability (Huang et al., 2022) and reduce training time without affecting test performance (Qin et al., 2024). Thus, we propose to incorporate the hardness of data points into the model optimization process when there is little to no label noise. Specifically, our optimization objective is to minimize $\frac{1}{N}\sum_{i\in N}h_{i}f(i;\theta)$ rather than usual empirical risk $\frac{1}{N}\sum_{i\in N}f(i;\theta)$, where $h_{i}$ is equal to $f(i;\theta)$ but does not participate in model backpropagation. Here, $h_{i}$ can also be interpreted as an unnormalized probability of selecting data point $i$ based on its hardness. Then the gradient of $i$-th data changes from $x_{i}$ to $h_{i}x_{i}$, and then we define the Compound of Hardness and Gradient score (CHG score) of a data subset $S$ as $U(S):=\|\frac{1}{N}\sum_{i\in N}h_{i}\nabla f(i;\theta)\|_{2}^{2}-\|\frac{1}{N}\sum_{i\in N}h_{i}\nabla f(i;\theta)-\frac{1}{S}\sum_{i\in S}h_{i}\nabla f(i;\theta)\|_{2}^{2}$. When the subset $S=\emptyset$, the produced utility $U(S)$ is 0.
However, when label noise cannot be ignored, the hardness metric becomes unreliable, as a well-performing classifier may assign a high hardness value to a training data point with an incorrect label. In this case, we still utilize the original score $\|\frac{1}{N}\sum_{i\in N}\nabla f(i;\theta)\|_{2}^{2}-\|\frac{1}{N}\sum_{i\in N}\nabla f(i;\theta)-\frac{1}{S}\sum_{i\in S}\nabla f(i;\theta)\|_{2}^{2},$ the corresponding method for calculating the Shapley value based on Equation 4 is referred to as GradE Shapley (Gradient Euclidean Shapley). We also denote the direct use of the hardness $h_{i}$ as the Shapley value of data point $i$ as Hardness Shapley. We compared these three methods across various dataset settings in our data selection experiments.
@@
Data Shapley (Ghorbani & Zou, 2019) Marginal contribution $\mathcal{O}(n^{2}\log n)$ + + ++ ++
@@
Beta Shapley (Kwon & Zou, 2022a) Marginal contribution $\mathcal{O}(n^{2}\log n)$ + + ++ ++
Data Banzhaf (Wang & Jia, 2023) Marginal contribution $\mathcal{O}(n\log n)$ + +
@@
Influence Function (Koh & Liang, 2017) Gradient NA - - + +
@@
Data-OOB (Kwon & Zou, 2023) Out-of-bag estimate NA ++ + - ++
CHG Shapley (this paper) Marginal contribution, $1$ ++ ++ ++ NA
@@
Table 1: A taxonomy of data valuation algorithms, mainly abstracted from (Jiang et al., 2023). The symbol “NA” means the method is not based on model training, then “No Answer”. The symbols ‘- / + / ++’ indicate that a corresponding data valuation method achieves a ‘similar or worse / better / much better’ performance than a random baseline, respectively. Detailed experimental results comparing these data valuation methods are provided in Appendix A.
3.3 CHG Shapley-based Data Valuation Algorithm
After obtaining the CHG score, which approximates the influence of data subsets on training loss using gradient and loss information during model training, we can derive the analytical expression for the Shapley value of each data point under this utility function. This method, denoted as CHG Shapley, enables efficient computation of the value of each data point in large-scale datasets.